22west:

help i only have 10 mins

22west:

1 attachment
EvieSwan2405:

Really? Your trying to get an answer and you told Ultri I gave you an answer. When you obviously want an answer because you said you had 10 minutes like a few minutes ago...

umm:

Restrain from arguing/fighting please?

22west:

hey you gave me a direct answer

22west:

and that goes against the qc rules

dude:

\[1\:\frac{1}{4}x=3\:\frac{1}{2}\] Solve for x

22west:

idk how

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude You can convert both equations into improper fractions \(\frac{5}{4}x=\frac{7}{2}\) And then multiply by the reciprocal of \(\frac{5}{4}\)(Which is \(\frac{4}{5}\)) \(\color{#0cbb34}{\text{End of Quote}}\)

22west:

ok

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude \(\huge{\color{red}{\frac{4}{5}\cdot} \frac{5}{4}x=\frac{7}{2}\color{red}{\cdot \frac{4}{5}}}\) \(\color{#0cbb34}{\text{End of Quote}}\)

22west:

ok

dude:

Now simplify that

22west:

how

dude:

Multiply both sides by \(\frac45\). Which is in red

22west:

1 and \[2\frac{ 4 }{ 5 }\]

dude:

That is your answer

22west:

which one

22west:

help?

22west:

which one is the answer